You are evaluating a project for The Farstroke golf club, guaranteed to correct that nasty slice. You estimate the sales price of The Farstroke to be $450 per unit and sales volume to be 1,200 units in year 1; 1,325 units in year 2; and 1,000 units in year 3. The project has a 3-year life. Variable costs amount to $250 per unit and fixed costs are $100,000 per year. The project requires an initial investment of $150,000 in assets, which will be depreciated straight-line to zero over the three-year project life. The actual market value of these assets at the end of year 3 is expected to be $30,000. NWC requirements at the beginning of each year will be approximately 20 percent of the projected sales during the coming year. The tax rate is 21 percent and the required return on the project is 10 percent.

What is the operating cash flow for the project in year 2?
Answer #1

1$ 596,250 =1325*450 =1325*250+100000 =150000/3 Sales Less: Costs Depreciation EBIT Less: Tax @ 21% Net income Add: Depreciat

*Please rate thumbs up